# Discrete Time Crystal Phase as a Resource for Quantum Enhanced Sensing

### v2

@@ Abstract
Conclusion.– Having established a DTC with indefinite persistent oscillations and strong robustness to the imperfections in the driving pulse, we show that this DTC is extremely precise in measuring coupling strength. It provide quantum-enhanced sensing over a region that extends over the DTC phase to the transition point. Through establishing finite-size scaling analysis, we characterize the nature of the phase transition as the second-order and determine relevant critical exponents. The obtained quantum-enhanced sensitivity is general and independent of the initial state. Regarding the imperfection in the rotating pulse, we show that increasing imperfection, before melting the DTC, enhances the precision of estimation in both DTC phase and transition point.
@@
Acknowledgement.– We thank Fernando Iemini for insightful comments. A.B. acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). R. Y. acknowledges support from the National Science Foundation of China for the International Young Scientists Fund (Grant No. 12250410242). This research was funded by the National Science Centre, Poland, Projects No. 2021/42/A/ST2/ 00017 (K.S.).
Supplementary Materials
I Robustness of the DTC against nonuniform imperfections
@@
$e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L/2}(-1)^{2}|\mathbf{5}\rangle$ (S1)
$e^{-iH_{p}}e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L/2}\Big{\{}c_{1}c_{2}c_{3}c_{4}|\mathbf{5}\rangle+(-i)c_{1}c_{2}c_{3}s_{4}|\mathbf{4}\rangle+(-i)c_{1}c_{2}s_{3}c_{4}|\mathbf{7}\rangle+(-i)^{2}c_{1}c_{2}s_{3}s_{4}|\mathbf{6}\rangle$ (S2)
$+$ $(-i)c_{1}s_{2}c_{3}c_{4}|\mathbf{1}\rangle+(-i)^{2}c_{1}s_{2}c_{3}s_{4}|\mathbf{0}\rangle+(-i)^{2}c_{1}s_{2}s_{3}c_{4}|\mathbf{3}\rangle+(-i)^{3}c_{1}s_{2}s_{3}s_{4}|\mathbf{2}\rangle$ (S3)
$+$ $(-i)s_{1}c_{2}c_{3}c_{4}|\mathbf{13}\rangle+(-i)^{2}s_{1}c_{2}c_{3}s_{4}|\mathbf{12}\rangle+(-i)^{2}s_{1}c_{2}s_{3}c_{4}|\mathbf{15}\rangle+(-i)^{3}s_{1}c_{2}s_{3}s_{4}|\mathbf{14}\rangle$ (S4)
$+$ $(-i)^{2}s_{1}s_{2}c_{3}c_{4}|\mathbf{9}\rangle+(-i)^{3}s_{1}s_{2}c_{3}s_{4}|\mathbf{8}\rangle+(-i)^{3}s_{1}s_{2}s_{3}c_{4}|\mathbf{11}\rangle+(-i)^{4}s_{1}s_{2}s_{3}s_{4}|\mathbf{10}\rangle\Big{\}}$ (S5)
$e^{-i\Omega H_{I}}e^{-iH_{p}}e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L}\Big{\{}(-1)^{2}c_{1}c_{2}c_{3}c_{4}|\mathbf{5}\rangle+(-1)^{1}(-i)c_{1}c_{2}c_{3}s_{4}|\mathbf{4}\rangle+(-1)^{3}(-i)c_{1}c_{2}s_{3}c_{4}|\mathbf{7}\rangle+(-1)^{2}(-i)^{2}c_{1}c_{2}s_{3}s_{4}|\mathbf{6}\rangle$ (S7)
$+$ $(-1)^{1}(-i)c_{1}s_{2}c_{3}c_{4}|\mathbf{1}\rangle+(-1)^{0}(-i)^{2}c_{1}s_{2}c_{3}s_{4}|\mathbf{0}\rangle+(-1)^{2}(-i)^{2}c_{1}s_{2}s_{3}c_{4}|\mathbf{3}\rangle+(-1)^{1}(-i)^{3}c_{1}s_{2}s_{3}s_{4}|\mathbf{2}\rangle$ (S8)
$+$ $(-1)^{3}(-i)s_{1}c_{2}c_{3}c_{4}|\mathbf{13}\rangle+(-1)^{2}(-i)^{2}s_{1}c_{2}c_{3}s_{4}|\mathbf{12}\rangle+(-1)^{4}(-i)^{2}s_{1}c_{2}s_{3}c_{4}|\mathbf{15}\rangle+(-1)^{3}(-i)^{3}s_{1}c_{2}s_{3}s_{4}|\mathbf{14}\rangle$ (S9)
$+$ $(-1)^{2}(-i)^{2}s_{1}s_{2}c_{3}c_{4}|\mathbf{9}\rangle+(-1)^{1}(-i)^{3}s_{1}s_{2}c_{3}s_{4}|\mathbf{8}\rangle+(-1)^{3}(-i)^{3}s_{1}s_{2}s_{3}c_{4}|\mathbf{11}\rangle+(-1)^{2}(-i)^{4}s_{1}s_{2}s_{3}s_{4}|\mathbf{10}\rangle\Big{\}}$ (S10)
$e^{-iH_{p}}e^{-i\Omega H_{I}}e^{-iH_{p}}e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L}\Big{\{}(c_{1}c_{2}c_{3}c_{4})^{2}+(c_{1}c_{2}c_{3}s_{4})^{2}+(c_{1}c_{2}s_{3}c_{4})^{2}+(c_{1}c_{2}s_{3}s_{4})^{2}+(c_{1}s_{2}c_{3}c_{4})^{2}+(c_{1}s_{2}c_{3}s_{4})^{2}$ (S12)
$+$ $(c_{1}s_{2}s_{3}c_{4})^{2}+(c_{1}s_{2}s_{3}s_{4})^{2}+(s_{1}c_{2}c_{3}c_{4})^{2}+(s_{1}c_{2}c_{3}s_{4})^{2}+(s_{1}c_{2}s_{3}c_{4})^{2}+(s_{1}c_{2}s_{3}s_{4})^{2}$ (S13)
$+$ $(s_{1}s_{2}c_{3}c_{4})^{2}+(s_{1}s_{2}c_{3}s_{4})^{2}+(s_{1}s_{2}s_{3}c_{4})^{2}+(s_{1}s_{2}s_{3}s_{4})^{2}\Big{\}}|\mathbf{5}\rangle+\cdots$ (S14)
$e^{-iH_{p}}e^{-i\Omega H_{I}}e^{-iH_{p}}e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L}|\mathbf{5}\rangle+\cdots$ (S16)
@@
II Role of initial state and imperfection effect
@@
Figure S1: (a)-(c) The QFI after $n{=}100$ period cycles as a function of deviation $\omega$ obtained for different values of $\varepsilon{\in}\{0.02,0.04,\cdots,0.3\}$ in a chain of size $L{=}12$ that is initialized in $|\psi_{0}\rangle{=}|\mathbf{0}\rangle{=}|{\uparrow,\uparrow,\cdots,\uparrow}\rangle$, a random state, and Néel state $|\psi_{0}\rangle{=}|\mathbf{1365}\rangle{=}|{\uparrow,\downarrow,\cdots,\uparrow,\downarrow}\rangle$, respectively.
@@
In the main text, we present results only for the initial state $|\psi_{0}\rangle{=}|\mathbf{0}\rangle{=}|{\uparrow,\cdots,\uparrow}\rangle$. In this section, we provide results for other computational states to show the generality of the observed behavior concerning $\omega$ and $\varepsilon$. In Fig. S1, we depict the QFI $\mathcal{F}_{Q}(nT)$ as a function of $\omega$ after $n{=}100$ period cycles in a system of size $L{=}12$, initialized in (a) $|\psi_{0}\rangle{=}|{\uparrow,\cdots,\uparrow}\rangle$, (b) a random state, and (c) the Néel state $|\psi_{0}\rangle{=}|{\uparrow,\downarrow\cdots,\uparrow,\downarrow}\rangle$. Curves with different colors correspond to different values of imperfections $\varepsilon$. In terms of $\omega$, one can see that the distinctive behavior of our system in both DTC and non-DTC phases reflects itself in all the considered initial states. Regarding the imperfection effect, by increasing $\varepsilon$, more information about $\omega$ can be printed in the evolved state, resulting in higher values of the QFI. As is clear from Fig. S1, this behavior is qualitatively independent of the initial states.
III Melting transition of the DTC
@@
In the main text, we show how increasing imperfection in the rotating pulse $\varepsilon$ can onset a phase transition between stable DTC and non-DTC phase. The transition driven by $\varepsilon$ is diagnosed by averaged entanglement entropy $\langle S_{EE}\rangle$ and averaged diagonal entropy $\langle S_{DE}\rangle$ obtained for all Floquet states of $U_{F}(\Omega,\varepsilon){=}\sum_{k=1}^{2^{L}}e^{-i\phi_{k}}|\phi_{k}\rangle\langle\phi_{k}|$. In the DTC phase, each Floquet state $|\phi_{k}\rangle$ is a maximally entangled GHZ state of a pair of computational basis states. For instance for small values of $\varepsilon$ one approximately has $|\phi_{1}\rangle{\cong}\frac{1}{\sqrt{2}}(|\mathbf{0}\rangle+|\mathbf{2^{L}-1}\rangle)$ and $|\phi_{2^{L}}\rangle{\cong}\frac{1}{\sqrt{2}}(|\mathbf{0}\rangle-|\mathbf{2^{L}-1}\rangle)$ with the corresponding eigenvalues as $\phi_{1}{\cong}E_{\mathbf{0}}$ and $\phi_{2^{L}}{\cong}E_{\mathbf{0}}\pm\pi$. Note that $\{E_{\mathbf{z}}\}$ are the eigenvalues of $H_{I}$ with $E_{\mathbf{z}}{=}E_{2^{L}-1-\mathbf{z}}$. Clearly, deep inside the DTC regime, the entanglement entropy is $S_{EE}^{(k)}{\cong}\ln 2$ for all $k$’s. This can be seen in Fig. S2 (a)-(c) which depict the averaged entanglement entropy $\langle S_{EE}\rangle$ for systems of various sizes and $\omega\in\{1,3,5\}\times 10^{-2}$. In this regime, by enlarging $\varepsilon$ the averaged entanglement entropy gets distance from $\ln 2$ and peaks at its size- and $\omega$-dependent location, happening for $\varepsilon{=}0.5$. The behavior of the entanglement entropy concerning $L$ hints that the melting transition is of second-order type. This means that one can extract the critical properties for the transition by implementing finite-size scaling analysis. However, as $\omega_{\max}{\propto}L^{-1}$, by increasing the system size the range of $\omega$’s that the DTC phase is stable for them, namely $\omega{<}\omega_{\max}$, shrinks. Therefore, the results for finite-size scaling analysis obtained using probes that for any given $\omega{<}\omega_{\max}$ these systems are within the DTC phase when $\varepsilon{\simeq}0$. Here, the numerical restriction in the ED method limits us to the system up to $L{=}12$. Presuming that the averaged entanglement entropy follows an ansatz as $\langle S_{EE}\rangle=L^{\zeta/\nu}\mathcal{D}(L^{1/\nu}(\varepsilon-\varepsilon_{c}))$, then plotting $L^{-\zeta/\nu}\langle S_{EE}\rangle$ as a function of $L^{1/\nu}(\varepsilon-\varepsilon_{c})$ collapses the curves of different sizes. Here, $\zeta$ and $\nu$ are the critical exponents, $\varepsilon_{c}$ is the critical point, and $\mathcal{D}$ is an arbitrary function. The best data collapse can be obtained for the optimal critical parameters $(\varepsilon_{c},\zeta,\nu)$. In the insets of Fig. S2(a)-(c), we present the best data collapse of the corresponding curves obtained for the reported critical parameters. For the sake of completeness, we repeat the analysis above for the averaged diagonal entropy. The diagonal entropy contains partial information about the system by setting the off-diagonal terms of the half-system reduced density matrix to zero. In Fig. S2(d)-(f), we present the obtained $\langle S_{DE}\rangle$ for various sizes and $\omega$’s. Indeed, the behavior of the averaged diagonal entropy is qualitatively close to the averaged entanglement entropy in both DTC and non-DTC phases. In particular, for small values of $\varepsilon$, one has $S_{EE}^{(k)}=S_{DE}^{(k)}=\ln 2$. In the insets of Fig. S2(d)-(f), we present the results of the finite-size scaling analysis that we established for this quantity. Surprisingly the obtained $\varepsilon_{c}$ for both $\langle S_{EE}\rangle$ and $\langle S_{DE}\rangle$ are close. Note that in the transition between the DTC and non-DTC phases driven by $\varepsilon$, one may not observe quantum-enhanced sensitivity in the process of sensing $\varepsilon$.
@@
Figure S2: (a)-(c) the averaged entanglement entropy $\langle S_{EE}\rangle$ as a function of pulse imperfection $\varepsilon$ obtained for systems of different sizes and different $\omega$’s. (d)-(f) The averaged diagonal entropy $\langle S_{DE}\rangle$ as a function of $\varepsilon$ in systems with various size and $\omega$’s. The insets of the panels show the results of the finite-size scaling analysis and optimal data collapse of curves with different sizes, which happen for the reported critical properties $(\varepsilon_{c},\nu,\zeta)$.
IV Characterizing the non-DTC region
As has been discussed in the main text, by increasing the divination $\omega$, the perfect and stable revivals of the fidelity in stroboscopic times, namely $F(2nT){=}1$ for $\omega{\leq}\omega_{\max}$, are replaced by nontrivial oscillations for $\omega{>}\omega_{\max}$. This hints one enters a non-DTC region. In this section we aim to characterize the nature of this region. Our results for the revival fidelity $F(2nT)$ as a function of $n$ for systems of different sizes that are tuned to work in the non-DTC region, namely for $\omega{\simeq}0.15$, have been illustrated in Fig. S3(a). By enlarging the system size, the period of these incommensurate fluctuations increases. This implies that, in systems with enough large sizes, these oscillations practically vanish in a reasonable time window, signaling the thermalization of the system. This observation receives more support from our static study based on the entanglement entropy and diagonal entropy. In a thermal system, the Floquet states $\{|\phi_{k}\rangle\}$ are expected to behave as a typical random pure state, therefore their entanglement entropy is predicted to follow the Page entropy $\langle S^{P}_{EE}\rangle{\simeq}(L\ln(2)-1)/2$ for enough large $L$’s. In this case, the average entanglement entropy should already captured its maximum and the variations of $\varepsilon$ may not considerably affect $\langle S^{P}_{EE}\rangle$. Our numerical results in Fig. S3(b) support this prediction for $\omega{=}\frac{\pi}{4}$, namely deep inside the thermal phase. The results for systems of size $L{=}8$ and $L{=}12$ capture the Page entropy $\langle S_{EE}\rangle{\simeq}\langle S^{P}_{EE}\rangle$ (depicted by colored dashed lines) with slight changes in terms of $\varepsilon$. Regarding the diagonal entropy, typical random pure states are expected to follow $\langle S^{P}_{DE}\rangle{\simeq}\ln(0.48\times 2^{L/2})+\ln(2)$. The presented results in Fig. S3(c) confirm this behavior. In this panel $\langle S^{P}_{DE}\rangle$ is represented by colored dashed lines.
Figure S3: (a) the stroboscopic dynamics of the revival fidelity $F(2nT)$ over $100$ period cycles when systems with different sizes evolve in non-DTC region with $\omega{\simeq}0.15$ and $\varepsilon{=}0.01$. (b) the averaged entanglement entropy $\langle S_{EE}\rangle$ and (c) the averaged diagonal entropy $\langle S_{DE}\rangle$ as a function of pulse imperfection $\varepsilon$ obtained for systems of different sizes and $\omega{=}\pi/4$. Dashed lines in (b) and (c) determine the corresponding $\langle S^{P}_{EE}\rangle$ and $\langle S^{P}_{DE}\rangle$, respectively.

### v3

@@ 9 Conclusion
We have established a DTC with indefinite persistent oscillations and strong robustness against imperfections in the driving pulse. We show that this DTC has an application for measuring the coupling strength with quantum-enhanced precision over a region that extends over the entire DTC phase. Through finite-size scaling analysis, we characterize the nature of the phase transition as the second-order and determine relevant critical exponents. In addition, we show that imperfection in the pulse, enhances the precision before melting the DTC. The proposed mechanism is independent of the initial state and can be realized in existing quantum simulators. This provides novel application of DTC phases in the field of quantum metrology.
10 Acknowledgement
@@
We thank Fernando Iemini for insightful comments. A.B. acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). R. Y. acknowledges support from the National Science Foundation of China for the International Young Scientists Fund (Grant No. 12250410242). This research was funded by the National Science Centre, Poland, Projects No. 2021/42/A/ST2/ 00017 (K.S.).
Appendix A Robustness of the DTC against nonuniform imperfections
@@
$e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L/2}(-1)^{2}|\mathbf{5}\rangle$ (A1)
$e^{-iH_{p}}e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L/2}\Big{\{}c_{1}c_{2}c_{3}c_{4}|\mathbf{5}\rangle+(-i)c_{1}c_{2}c_{3}s_{4}|\mathbf{4}\rangle+(-i)c_{1}c_{2}s_{3}c_{4}|\mathbf{7}\rangle$ (A2)
$+$ $(-i)^{2}c_{1}c_{2}s_{3}s_{4}|\mathbf{6}\rangle+(-i)c_{1}s_{2}c_{3}c_{4}|\mathbf{1}\rangle+(-i)^{2}c_{1}s_{2}c_{3}s_{4}|\mathbf{0}\rangle$ (A3)
$+$ $(-i)^{2}c_{1}s_{2}s_{3}c_{4}|\mathbf{3}\rangle+(-i)^{3}c_{1}s_{2}s_{3}s_{4}|\mathbf{2}\rangle+(-i)s_{1}c_{2}c_{3}c_{4}|\mathbf{13}\rangle$ (A4)
$+$ $(-i)^{2}s_{1}c_{2}c_{3}s_{4}|\mathbf{12}\rangle+(-i)^{2}s_{1}c_{2}s_{3}c_{4}|\mathbf{15}\rangle+(-i)^{3}s_{1}c_{2}s_{3}s_{4}|\mathbf{14}\rangle$ (A5)
$+$ $(-i)^{2}s_{1}s_{2}c_{3}c_{4}|\mathbf{9}\rangle+(-i)^{3}s_{1}s_{2}c_{3}s_{4}|\mathbf{8}\rangle+(-i)^{3}s_{1}s_{2}s_{3}c_{4}|\mathbf{11}\rangle$ (A6)
$+$ $(-i)^{4}s_{1}s_{2}s_{3}s_{4}|\mathbf{10}\rangle\Big{\}}$ (A7)
$e^{-i\Omega H_{I}}e^{-iH_{p}}e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L}\Big{\{}(-1)^{2}c_{1}c_{2}c_{3}c_{4}|\mathbf{5}\rangle+(-1)^{1}(-i)c_{1}c_{2}c_{3}s_{4}|\mathbf{4}\rangle$ (A8)
$+$ $(-1)^{3}(-i)c_{1}c_{2}s_{3}c_{4}|\mathbf{7}\rangle+(-1)^{2}(-i)^{2}c_{1}c_{2}s_{3}s_{4}|\mathbf{6}\rangle$ (A9)
$+$ $(-1)^{1}(-i)c_{1}s_{2}c_{3}c_{4}|\mathbf{1}\rangle+(-1)^{0}(-i)^{2}c_{1}s_{2}c_{3}s_{4}|\mathbf{0}\rangle$ (A10)
$+$ $(-1)^{2}(-i)^{2}c_{1}s_{2}s_{3}c_{4}|\mathbf{3}\rangle+(-1)^{1}(-i)^{3}c_{1}s_{2}s_{3}s_{4}|\mathbf{2}\rangle$ (A11)
$+$ $(-1)^{3}(-i)s_{1}c_{2}c_{3}c_{4}|\mathbf{13}\rangle+(-1)^{2}(-i)^{2}s_{1}c_{2}c_{3}s_{4}|\mathbf{12}\rangle$ (A12)
$+$ $(-1)^{4}(-i)^{2}s_{1}c_{2}s_{3}c_{4}|\mathbf{15}\rangle+(-1)^{3}(-i)^{3}s_{1}c_{2}s_{3}s_{4}|\mathbf{14}\rangle$ (A13)
$+$ $(-1)^{2}(-i)^{2}s_{1}s_{2}c_{3}c_{4}|\mathbf{9}\rangle+(-1)^{1}(-i)^{3}s_{1}s_{2}c_{3}s_{4}|\mathbf{8}\rangle$ (A14)
$+$ $(-1)^{3}(-i)^{3}s_{1}s_{2}s_{3}c_{4}|\mathbf{11}\rangle+(-1)^{2}(-i)^{4}s_{1}s_{2}s_{3}s_{4}|\mathbf{10}\rangle\Big{\}}$ (A15)
$e^{-iH_{p}}e^{-i\Omega H_{I}}e^{-iH_{p}}e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L}\Big{\{}(c_{1}c_{2}c_{3}c_{4})^{2}+(c_{1}c_{2}c_{3}s_{4})^{2}+(c_{1}c_{2}s_{3}c_{4})^{2}+(c_{1}c_{2}s_{3}s_{4})^{2}+(c_{1}s_{2}c_{3}c_{4})^{2}$ (A16)
$+$ $(c_{1}s_{2}c_{3}s_{4})^{2}+(c_{1}s_{2}s_{3}c_{4})^{2}+(c_{1}s_{2}s_{3}s_{4})^{2}+(s_{1}c_{2}c_{3}c_{4})^{2}$ (A17)
$+$ $(s_{1}c_{2}c_{3}s_{4})^{2}+(s_{1}c_{2}s_{3}c_{4})^{2}+(s_{1}c_{2}s_{3}s_{4})^{2}+(s_{1}s_{2}c_{3}c_{4})^{2}$ (A18)
$+$ $(s_{1}s_{2}c_{3}s_{4})^{2}+(s_{1}s_{2}s_{3}c_{4})^{2}+(s_{1}s_{2}s_{3}s_{4})^{2}\Big{\}}|\mathbf{5}\rangle+\cdots$ (A19)
$e^{-iH_{p}}e^{-i\Omega H_{I}}e^{-iH_{p}}e^{-i\Omega H_{I}}|\mathbf{5}\rangle=$ $(-i)^{L}|\mathbf{5}\rangle+\cdots$ (A20)
@@
Appendix B Imperfection effect and role of the initial state
Figure 1: (a) The QFI versus $\omega$ for different $\varepsilon$. The dashed line determines the onset of the phase transition. (b) The QFI versus $\varepsilon$ for two values of $\omega$. The results are obtained after $n{=}100$ cycling periods in a system of size $L{=}12$, initialized in $|\psi_{0}\rangle{=}|\mathbf{0}\rangle{=}|{\uparrow,\uparrow,\cdots,\uparrow}\rangle$.
@@ Appendix B Imperfection effect and role of the initial state
Figure 2: (a)-(b) The QFI after $n{=}100$ period cycles as a function of deviation $\omega$ obtained for different values of $\varepsilon{\in}\{0.02,0.04,\cdots,0.3\}$ in a chain of size $L{=}12$ that is initialized in Néel state $|\psi_{0}\rangle{=}|\mathbf{1365}\rangle{=}|{\uparrow,\downarrow,\cdots,\uparrow,\downarrow}\rangle$, and a random state, respectively.
In the main text, we analytically proved that our DTC is robust against uniform imperfection $\varepsilon$ in the driving pulse when $\omega{=}0$. In the case of nonzero $\omega$, the situation becomes even more interesting. In Fig. 1(a), we plot the QFI versus $\omega$ under driving pulse with various imperfections $\varepsilon$. While the qualitative behavior of the probe in the DTC phase is not affected by imperfection, increasing $\varepsilon$ enhances the QFI. This can be understood as imperfect rotating pulses through involving a larger sector of the Hilbert space in the dynamics of the system imprints more information about $\omega$ into the quantum state. Notably, the onset of the transition from DTC to non-DTC region, which reflects a peak in $\mathcal{F}_{Q}$, is almost independent of the imperfection value. To assess the performance in a wider range of the imperfection, in Fig. 1(b), $\mathcal{F}_{Q}$ as a function of $\varepsilon$ for deep inside the DTC phase, i.e. $\omega{=}10^{-4}\pi/2$, and at the transition point, i.e. $\omega{=}\omega_{\max}$, have been reported. The enhancement in the DTC phase, where the system is supposed to be strongly localized, has a remarkably stronger effect than at the transition point which already has features of both thermalization and localization. This interesting result is in sharp contrast with the usual sensors where the imperfections deteriorate the sensing power. The results are obtained after $n{=}100$ cycling period for a chain of size $L{=}12$ initialized in $|\psi_{0}\rangle{=}|\mathbf{0}\rangle{=}|{\uparrow,\uparrow,\cdots,\uparrow}\rangle$.
@@
As we mention in the main text, the observed behavior of the DTC sensor concerning $\omega$ and $\varepsilon$ is general and independent of the initial states. To support this claim, in Figs. 2 (a)-(b), we depict the QFI $\mathcal{F}_{Q}(nT)$ as a function of $\omega$ in a system initialized in the Néel state $|\psi_{0}\rangle{=}|{\uparrow,\downarrow\cdots,\uparrow,\downarrow}\rangle$ and a random state, respectively. The results are obtained after $n{=}100$ period cycles in a system of size $L{=}12$. Curves with different colors correspond to different values of imperfections $\varepsilon\in\{0.02,\cdots,0.3\}$. In terms of $\omega$, one can see that the distinctive behavior of our system in both DTC and non-DTC phases, as well as the transition between them at $\omega{=}\omega_{\max}$ reflects itself in all the considered initial states. Regarding the imperfection effect, by increasing $\varepsilon$, more information about $\omega$ can be printed in the evolved state, resulting in higher values of the QFI. As is clear from Fig. 2, this behavior is qualitatively independent of the initial states.
Appendix C Melting transition of the DTC
@@
In the main text, we show how increasing imperfection in the rotating pulse $\varepsilon$ can onset a phase transition between stable DTC and non-DTC phase. The transition driven by $\varepsilon$ is diagnosed by averaged entanglement entropy $\langle S_{EE}\rangle$ and averaged diagonal entropy $\langle S_{DE}\rangle$ obtained for all Floquet states of $U_{F}(\Omega,\varepsilon){=}\sum_{k=1}^{2^{L}}e^{-i\phi_{k}}|\phi_{k}\rangle\langle\phi_{k}|$. In the DTC phase, each Floquet state $|\phi_{k}\rangle$ is a maximally entangled GHZ state of a pair of computational basis states. For instance for small values of $\varepsilon$ one approximately has $|\phi_{1}\rangle{\cong}\frac{1}{\sqrt{2}}(|\mathbf{0}\rangle+|\mathbf{2^{L}-1}\rangle)$ and $|\phi_{2^{L}}\rangle{\cong}\frac{1}{\sqrt{2}}(|\mathbf{0}\rangle-|\mathbf{2^{L}-1}\rangle)$ with the corresponding eigenvalues as $\phi_{1}{\cong}E_{\mathbf{0}}$ and $\phi_{2^{L}}{\cong}E_{\mathbf{0}}\pm\pi$. Note that $\{E_{\mathbf{z}}\}$ are the eigenvalues of $H_{I}$ with $E_{\mathbf{z}}{=}E_{2^{L}-1-\mathbf{z}}$. Clearly, deep inside the DTC regime, the entanglement entropy is $S_{EE}^{(k)}{\cong}\ln 2$ for all $k$’s. This can be seen in Fig. 3 (a)-(c) which depict the averaged entanglement entropy $\langle S_{EE}\rangle$ for systems of various sizes and $\omega\in\{1,3,5\}\times 10^{-2}$. In this regime, by enlarging $\varepsilon$ the averaged entanglement entropy gets distance from $\ln 2$ and peaks at its size- and $\omega$-dependent location, happening for $\varepsilon{=}0.5$. The behavior of the entanglement entropy concerning $L$ hints that the melting transition is of second-order type. This means that one can extract the critical properties for the transition by implementing finite-size scaling analysis. However, as $\omega_{\max}{\propto}L^{-1}$, by increasing the system size the range of $\omega$’s that the DTC phase is stable for them, namely $\omega{<}\omega_{\max}$, shrinks. Therefore, the results for finite-size scaling analysis obtained using probes that for any given $\omega{<}\omega_{\max}$ these systems are within the DTC phase when $\varepsilon{\simeq}0$. Here, the numerical restriction in the ED method limits us to the system up to $L{=}12$. Presuming that the averaged entanglement entropy follows an ansatz as $\langle S_{EE}\rangle=L^{\zeta/\nu}\mathcal{D}(L^{1/\nu}(\varepsilon-\varepsilon_{c}))$, then plotting $L^{-\zeta/\nu}\langle S_{EE}\rangle$ as a function of $L^{1/\nu}(\varepsilon-\varepsilon_{c})$ collapses the curves of different sizes. Here, $\zeta$ and $\nu$ are the critical exponents, $\varepsilon_{c}$ is the critical point, and $\mathcal{D}$ is an arbitrary function. The best data collapse can be obtained for the optimal critical parameters $(\varepsilon_{c},\zeta,\nu)$.
@@
Figure 3: (a)-(c) the averaged entanglement entropy $\langle S_{EE}\rangle$ as a function of pulse imperfection $\varepsilon$ obtained for systems of different sizes and different $\omega$’s. (d)-(f) The averaged diagonal entropy $\langle S_{DE}\rangle$ as a function of $\varepsilon$ in systems with various size and $\omega$’s. The insets of the panels show the results of the finite-size scaling analysis and optimal data collapse of curves with different sizes, which happen for the reported critical properties $(\varepsilon_{c},\nu,\zeta)$.
In the insets of Fig. 3(a)-(c), we present the best data collapse of the corresponding curves obtained for the reported critical parameters. For the sake of completeness, we repeat the analysis above for the averaged diagonal entropy. The diagonal entropy contains partial information about the system by setting the off-diagonal terms of the half-system reduced density matrix to zero. In Fig. 3(d)-(f), we present the obtained $\langle S_{DE}\rangle$ for various sizes and $\omega$’s. Indeed, the behavior of the averaged diagonal entropy is qualitatively close to the averaged entanglement entropy in both DTC and non-DTC phases. In particular, for small values of $\varepsilon$, one has $S_{EE}^{(k)}=S_{DE}^{(k)}=\ln 2$. In the insets of Fig. 3(d)-(f), we present the results of the finite-size scaling analysis that we established for this quantity. Surprisingly the obtained $\varepsilon_{c}$ for both $\langle S_{EE}\rangle$ and $\langle S_{DE}\rangle$ are close. Note that in the transition between the DTC and non-DTC phases driven by $\varepsilon$, one may not observe quantum-enhanced sensitivity in the process of sensing $\varepsilon$.